# Scaling Beyond Context: A Survey of Multimodal Retrieval-Augmented Generation for Document Understanding

### v1

@@ Abstract
Document understanding is critical for applications from financial analysis to scientific discovery. Current approaches, whether OCR-based pipelines feeding Large Language Models (LLMs) or native Multimodal LLMs (MLLMs), face key limitations: the former loses structural detail, while the latter struggles with context modeling. Retrieval-Augmented Generation (RAG) helps ground models in external data, but documents’ multimodal nature, i.e., combining text, tables, charts, and layout, demands a more advanced paradigm: Multimodal RAG. This approach enables holistic retrieval and reasoning across all modalities, unlocking comprehensive document intelligence. Recognizing its importance, this paper presents a systematic survey of Multimodal RAG for document understanding. We propose a taxonomy based on domain, retrieval modality, and granularity, and review advances involving graph structures and agentic frameworks. We also summarize key datasets, benchmarks, and applications, and highlight open challenges in efficiency, fine-grained representation, and robustness, providing a roadmap for future progress in document AI.
@@ 1 Introduction
Document understanding has become a pivotal task in the era of information explosion, as it empowers machines to automatically interpret, organize, and reason over the massive volumes of unstructured and semi-structured documents produced across diverse domains (Subramani et al., 2020; Ding et al., 2024). Early studies primarily focused on text-centric documents, relying on optical character recognition (OCR) techniques (Gu et al., 2021; Appalaraju et al., 2021; Shi et al., 2016) to support layout analysis and key information extraction. However, in real-world scenarios, particularly in scientific domains, documents are often visually rich and contain complex elements such as tables, charts, and images (Park et al., 2019; Ding et al., 2025a). With the rapid progress of Large Language Models (LLMs) and the rising demand for understanding increasingly complex and diverse document types, developing robust and generalizable document understanding frameworks has become an area of growing interest.
@@
In visually rich document understanding, different approaches have emerged to address the challenges of integrating layout, text, and structural information. Multimodal LLM (MLLM)-native methods commonly represent documents as long image sequences, enabling unified learning across modalities with MLLMs (Duan et al., 2025; Xiong et al., 2025; Yu et al., 2025b; Zhou et al., 2024; Nassar et al., 2025; Ye et al., 2023; Hu et al., 2024a, b). While effective, these models struggle with very long documents spanning hundreds or thousands of pages, where sequence length limitations can hinder accurate retrieval and increase the risk of hallucination (Deng et al., 2024; Ma et al., 2024b). To improve modularity and robustness, agent-based approaches introduce specialized agents for subtasks such as layout analysis, content extraction, instruction decomposition, and verification (Liu et al., 2025; Han et al., 2025; Wang et al., 2025a; Wu et al., ; Yu et al., 2025c), though such designs often increase system complexity due to coordination overhead. Retrieval-augmented generation (RAG) methods provide another direction by grounding responses with external knowledge, typically retrieving the top-K most relevant pages (see Figure 1 (a)) across one or more documents (Lewis et al., 2020). Importantly, these paradigms are not mutually exclusive: RAG-based systems may employ agents to manage retrieval and verification, while agent-based workflows often incorporate RAG as one of the agent nodes, yielding more flexible hybrid frameworks. These complementary perspectives have shaped the landscape of document understanding, yet among them, RAG has drawn particular attention for its practicality and rapid growth (Arslan et al., 2024; Fan et al., 2024).
@@
Early RAG studies mainly rely on text-centric strategies, extracting text via OCR or combining OCR with MLLM-generated captions for visually rich documents, followed by encoding for retrieval (Wang et al., 2022; Li et al., 2023; Chen et al., 2024c; Khattab and Zaharia, 2020). Despite their effectiveness in certain scenarios, such text-based approaches exhibit fundamental limitations in handling visually rich documents, as they fail to adequately capture cross-modal cues and structural semantics (Abootorabi et al., 2025; Mei et al., 2025). To address these shortcomings, recent efforts have increasingly focused on multimodal RAG frameworks. The growth trend in the number of papers is shown in Figure 1 (b). These methods often represent multi-page documents as image sequences (Faysse et al., 2024; Yu et al., 2024), enabling visual encoders to extract richer representations for retrieval. Recent advances in multimodal RAG have increasingly emphasized finer-grained modeling within individual pages, including tables, charts, and other structured elements, to improve retrieval accuracy and robustness (Wang et al., 2025c; Choi et al., 2025). Extending beyond these coarse-to-fine refinements, recent studies have also investigated graph-based indexing Yuan et al. (2025) and multi-agent frameworks (Liu et al., 2025), which provide complementary mechanisms for structured reasoning and collaborative coordination in multimodal RAG.
@@ 2 Preliminary
DSE (2024a) EMNLP Phi3V CLIP-ViT-L/14 Open Image Page
@@
CRAEM (2024a) ACM MM LLaMA2-7B Pix2Struct Closed Image+Text Page
@@
GME (2025b) CVPR Qwen2-VL-7B ViT-BigG Open Image+Text Page
@@
HM-RAG (2025) ACM MM Qwen2.5-VL-7B ViT-BigG Open Image+Text Page
@@
FRAG (2025) Preprint InternVL2-8B InternViT-300M Closed Image Page
@@
Patho-AgenticRAG (2025a) Preprint Qwen2-VL-2B ViT-BigG Open Image Page
@@
CMRAG (2025) Preprint Qwen2.5-VL-7B-Instruct ViT-BigG Open Image+Text Page
MoLoRAG (2025) EMNLP Qwen2.5-VL-7B ViT-BigG Open Image Page
SERVAL (2025) Preprint InternVL3-14B InternViT-300M Open Image Page
@@ 3.1 Open and Closed Domain
Open-Domain Multimodal RAG. Open-domain multimodal RAG enhances an LLM’s knowledge in specialized domains by constructing retrieval databases from large collections of documents. Early approaches typically apply OCR to all documents to build text-based retrieval indices (Wang et al., 2022; Li et al., 2023; Chen et al., 2024c; Khattab and Zaharia, 2020), but this process is computationally expensive and inefficient. To improve scalability, recent methods such as DSE (Ma et al., 2024a) and ColPali (Faysse et al., 2024) leverage vision-language models (VLMs) to encode document pages directly into image embeddings, achieving significant efficiency gains. Despite these advances, most approaches still focus on reasoning within single documents and lack explicit mechanisms for integrating knowledge across sources. Addressing this limitation, M3DocRAG (Cho et al., 2024a) introduces approximate indexing to accelerate large-scale retrieval and establishes the benchmark M3DoCVQA with over 3,000 documents, while VDocRAG (Tanaka et al., 2025) constructs the OpenDocVQA dataset and mitigates page-level information loss by compressing visual content into dense token representations aligned with text.
@@
Closed-Domain Multimodal RAG. Closed-domain multimodal RAG is designed for practical scenarios where MLLMs encounter difficulties with extremely long documents or videos. Current MLLMs remain constrained by limited context windows, and long-context processing often amplifies the risk of hallucination. To address this, closed-domain approaches retrieve only the most relevant segments (e.g., pages or frames) from a target document and provide them as input to the MLLM, thereby improving both efficiency and reliability. For single-document visual question answering (DocVQA), SV-RAG (Chen et al., 2024b) employs the MLLM itself as a multimodal retriever, with specialized adapters for page retrieval and evidence-based reasoning. FRAG (Huang et al., 2025), by contrast, independently scores each frame or page, applies a Top-K selection to retain the most informative content, and then delegates answer generation to existing LMMs. CREAM (Zhang et al., 2024a) introduces a coarse-to-fine multimodal retrieval and attention-pooling integration framework, enabling effective cross-page reasoning and multi-page document comprehension for visual question answering. All approaches demonstrate that closed-domain multimodal RAG enables effective comprehension of long documents and videos without extending the model’s context length.
@@ 3.2 Retrieval Modality
Image-based Retrieval Modality. To handle visually rich documents, most existing methods represent each page as an image and encode it with VLMs, using the VLMs’ hidden states as page-level representations (see Figure 3 (a)). In parallel, the query is encoded, and page–query relevance is computed via similarity-based ranking (Ma et al., 2024a; Faysse et al., 2024; Yu et al., 2024; Chen et al., 2024b; Ma et al., 2025; Yu et al., 2025a). Building on image-based embeddings, MM-R5 (Xu et al., 2025a) introduces a reasoning-enhanced reranker that combines supervised fine-tuning and reinforcement learning to strengthen instruction following, elicit explicit reasoning chains, and leverage task-specific rewards for greater precision and interpretability. Complementing this direction, Light-ColPali (Ma et al., 2025) targets memory-efficient visual document retrieval by reducing patch-level embeddings through token merging, achieving near-optimal retrieval with a fraction of the original memory footprint.
Image+Text based Retrieval Modality. Leveraging both image and text for retrieval mitigates the loss of fine-grained textual cues that arise when relying solely on page-level VLM encoders. The text channel is derived from OCR (Zhang et al., 2024a; Suri et al., 2025; Liu et al., 2025; Wang et al., 2025b) or from summary annotations generated by large VLMs (Jain et al., 2025; Choi et al., 2025) (see Figure 2 (b)). VisDomRAG (Suri et al., 2025) and HM-RAG (Liu et al., 2025) adopt a dual-path pipeline: they retrieve and reason within each modality, then summarize and fuse the results into a single answer. By contrast, ViDoRAG (Wang et al., 2025b) and PREMIR (Choi et al., 2025) also retrieve per modality but merge candidates via a simple union before answer generation. Complementing these designs, SimpleDoc (Jain et al., 2025) uses a two-stage scheme for DocVQA: embedding-based candidate selection followed by re-ranking with VLM-generated page summaries, so that the summaries provide richer semantics for more precise evidence aggregation.
@@ 3.3 Retrieval Granularity
VRAG-RL (Wang et al., 2025c) equips LLMs with grounding capabilities through reinforcement learning, enabling the model to focus on finer-grained regions within retrieved pages that directly relate to the query. In a complementary way, MG-RAG (Xu et al., 2025b) adopts a multi-granularity multimodal retrieval strategy with hierarchical encoding and layout-aware search, which allows retrieval not only at the page level but also within specific structures such as tables and images. Similarly, DocVQA-RAP (Yu et al., 2025a) introduces a utility-driven retrieval mechanism that prioritizes document segments based on their contribution to answer quality, thereby filtering out fine-grained but redundant evidence. Along this line, MMRAG-DocQA (Gong et al., 2025) employs hierarchical indexing and multi-granularity semantic retrieval to capture fine-grained multimodal associations and long-distance dependencies across pages. Furthermore, mKG-RAG (Yuan et al., 2025) leverages multimodal knowledge graphs and a dual-stage retrieval strategy to refine fine-grained evidence selection by aligning entities and relationships across modalities. Finally, PREMIR (Choi et al., 2025) conducts fine-grained retrieval inside pages by generating predefined QA pairs for tables and charts, and matching them against queries to enhance precision. Together, these approaches illustrate a clear trend toward increasingly fine-grained retrieval in multimodal and document-heavy RAG systems.
@@ 3.4 Hybrid Enhancements for Multimodal RAG
HM-RAG (Liu et al., 2025) introduces a hierarchical multi-agent framework where graph-based databases serve as one of the critical modalities for multi-source retrieval, enabling the system to capture structured relations that complement unstructured and web-based evidence. In contrast, mKG-RAG (Yuan et al., 2025) explicitly constructs multimodal knowledge graphs by aligning entities and relationships across vision and text, using them as structured repositories to refine retrieval precision and reliability in knowledge-intensive VQA tasks. Building on this idea, DB3Team-RAG (Xia et al., 2025) incorporates image-indexed knowledge graphs within its domain-specific retrieval pipelines, allowing the system to handle complex multi-turn and ego-centric queries with structured multimodal context. Distinct from these approaches, MoLoRAG (Wu et al., 2025) leverages a page graph that encodes logical connections between document pages, using graph traversal to retrieve semantically and logically relevant evidence for multi-page document understanding. Collectively, these methods underscore the central role of graph structures, whether as retrieval sources, structured repositories, or logical page representations, in advancing fine-grained and reliable retrieval within multimodal RAG systems.
@@
Agent-based Multimodal RAG
Agent-based multimodal RAG extends the paradigm by deploying autonomous agents that orchestrate retrieval–generation interactions across modalities. These agents dynamically formulate queries, choose retrieval strategies, and adaptively fuse information from text, images, tables, and other modalities according to task requirements (see Figure 5 (b)). Multi-agent collaboration further supports iterative reasoning, verification, and evidence refinement, thereby enhancing the accuracy, reliability, and transparency of multimodal RAG systems.
ViDoRAG (Wang et al., 2025b) introduces an iterative agent workflow in which agents are responsible for exploration, summarization, and reflection, thereby enhancing multimodal retrieval and reasoning across visually rich documents. In comparison, HM-RAG (Liu et al., 2025) designs a hierarchical multi-agent architecture composed of a Decomposition Agent for query rewriting and context augmentation, modality-specific Retrieval Agents for parallel evidence collection from diverse sources, and a Decision Agent that integrates results through consistency voting and refinement. Extending this agent-based paradigm into the medical domain, Patho-AgenticRAG (Zhang et al., 2025a) employs agents capable of task decomposition and multi-turn search interactions, enabling accurate retrieval of joint text–image evidence from pathology textbooks while mitigating hallucinations in diagnostic reasoning. Taken together, these frameworks exemplify how specialized agent designs, ranging from iterative reasoning to hierarchical orchestration and domain-tailored search, advance the fine-grained retrieval and reasoning capabilities of multimodal RAG systems.

### v2

@@ Abstract
Document understanding is critical for applications from financial analysis to scientific discovery. Current approaches, whether OCR-based pipelines feeding Large Language Models (LLMs) or native Multimodal LLMs (MLLMs), face key limitations: the former loses structural detail, while the latter struggles with context modeling. Retrieval-Augmented Generation (RAG) helps ground models in external data, but documents’ multimodal nature, i.e., combining text, tables, charts, and layout, demands a more advanced paradigm: Multimodal RAG. This approach enables holistic retrieval and reasoning across all modalities, unlocking comprehensive document intelligence. Recognizing its importance, this paper presents a systematic survey of Multimodal RAG for document understanding. We propose a taxonomy based on domain, retrieval modality, and granularity, and review advances involving graph structures and agentic frameworks. We also summarize key datasets, benchmarks, applications and industry deployment, and highlight open challenges in efficiency, fine-grained representation, and robustness, providing a roadmap for future progress in document AI.
@@ 1 Introduction
Document understanding has become a pivotal task in the era of information explosion, as it empowers machines to automatically interpret, organize, and reason over the massive volumes of unstructured and semi-structured documents produced across diverse domains (Subramani et al., 2020; Ding et al., 2024). Early studies primarily focus on text-centric documents, relying on optical character recognition (OCR) techniques (Gu et al., 2021; Appalaraju et al., 2021; Shi et al., 2016) to support layout analysis and key information extraction. However, in real-world scenarios, particularly in scientific domains, documents are often visually rich and contain complex elements such as tables, charts, and images (Park et al., 2019; Ding et al., 2025a). With the rapid progress of Large Language Models (LLMs) and the rising demand for understanding increasingly complex and diverse document types, developing robust and generalizable document understanding frameworks has become an area of growing interest.
Benchmark Scope # Pages Visual Tokens
M3DocVQA (Cho et al., 2024a) Open-Domain $\sim$40K $\sim$41M
VisDoMBench (Suri et al., 2025) Open-Domain $\sim$21K $\sim$21M
OpenDocVQA (Tanaka et al., 2025) Open-Domain $\sim$206K $\sim$206M
Table 1: Scale of representative document RAG benchmarks. Visual tokens are estimated assuming $\sim$1K visual tokens per page.
@@
In visually rich document understanding, different approaches have emerged to address the challenges of integrating layout, text, and structural information. Multimodal LLM (MLLM)-native methods commonly represent documents as long image sequences, enabling unified learning across modalities with MLLMs (Duan et al., 2025; Xiong et al., 2025; Yu et al., 2025c; Zhou et al., 2024; Nassar et al., 2025; Ye et al., 2023; Hu et al., 2024a, b). While effective, these models struggle with very long documents spanning hundreds or thousands of pages, where sequence length limitations can hinder accurate retrieval and increase the risk of hallucination (Deng et al., 2024a; Ma et al., 2024c). As shown in Table 1, current multimodal RAG benchmarks require 20–200M visual tokens, far exceeding the typical 128K–1M context limits of existing MLLMs (Yang et al., 2025; Achiam et al., 2023; Team et al., 2023). To improve modularity and robustness, agent-based approaches introduce specialized agents for subtasks such as layout analysis, content extraction, instruction decomposition, and verification (Liu et al., 2025b; Han et al., 2025; Wang et al., 2025a; Wu et al., ; Yu et al., 2025d), though such designs often increase system complexity due to coordination overhead. Retrieval-augmented generation (RAG) methods provide another direction by grounding responses with external knowledge, typically retrieving the top-K most relevant pages (see Figure 1 (a)) across one or more documents (Lewis et al., 2020). Importantly, these paradigms are not mutually exclusive: RAG-based systems may employ agents to manage retrieval and verification, while agent-based workflows often incorporate RAG as one of the agent nodes, yielding more flexible hybrid frameworks. These complementary perspectives have shaped the landscape of document understanding, yet among them, RAG has drawn particular attention for its practicality and rapid growth (Arslan et al., 2024; Fan et al., 2024).
@@
Early RAG studies mainly rely on text-centric strategies, extracting text via OCR or combining OCR with MLLM-generated captions for visually rich documents, followed by encoding for retrieval (Wang et al., 2022; Li et al., 2023; Chen et al., 2024c; Khattab and Zaharia, 2020). Despite their effectiveness in certain scenarios, such text-based approaches exhibit fundamental limitations in handling visually rich documents, as they fail to adequately capture cross-modal cues and structural semantics (Abootorabi et al., 2025; Mei et al., 2025). To address these shortcomings, recent efforts have increasingly focused on multimodal RAG frameworks. The growth trend in the number of papers is shown in Figure 1 (b). These methods often represent multi-page documents as image sequences (Faysse et al., 2024; Yu et al., 2024), enabling visual encoders to extract richer representations for retrieval. Recent advances in multimodal RAG have increasingly emphasized finer-grained modeling within individual pages, including tables, charts, and other structured elements, to improve retrieval accuracy and robustness (Wang et al., 2025c; Choi et al., 2025). Extending beyond these coarse-to-fine refinements, recent studies have also investigated graph-based indexing Yuan et al. (2025) and multi-agent frameworks (Liu et al., 2025b), which provide complementary mechanisms for structured reasoning and collaborative coordination in multimodal RAG.
@@ 2 Preliminary
DSE (2024b) EMNLP Phi3V CLIP-ViT-L/14 Open Image Page
@@
CREAM (2024a) ACM MM LLaMA2-7B Pix2Struct Closed Image+Text Page
@@
GME (2025d) CVPR Qwen2-VL-7B ViT-BigG Open Image+Text Page
@@
HM-RAG (2025b) ACM MM Qwen2.5-VL-7B ViT-BigG Open Image+Text Page
@@
FRAG (2025a) Preprint InternVL2-8B InternViT-300M Closed Image Page
@@
Patho-AgenticRAG (2025c) Preprint Qwen2-VL-2B ViT-BigG Open Image Page
@@
CMRAG (2025c) Preprint Qwen2.5-VL-7B-Instruct ViT-BigG Open Image+Text Page
MoLoRAG (2025b) EMNLP Qwen2.5-VL-7B ViT-BigG Open Image Page
SERVAL (2025b) Preprint InternVL3-14B InternViT-300M Open Image Page
@@
RECON (Wang and Chen, 2025) Preprint GPT-4o-mini Open Image+Text Element
LAD-RAG (Sourati et al., 2025) Preprint GPT-4o-200b-128 Open Image+Text Element
HEAVEN (Kim et al., 2025) Preprint Qwen2.5-VL-3B-Instruct ViT-BigG Open Image Page
DREAM (Zhang et al., 2025a) ACM MM InternVL2-40B InternViT-6B Closed Image Page
MARA (Wu et al., 2025a) ACM MM MiniCPM-V2.0 SigLIP-SO400M Open Image Element
HEAR (Chen et al., 2025a) ACM MMW Qwen2.5-VL-32B-Instruct ViT-BigG Closed Image+Text Page
HPC-ColPali (Bach, 2025) Preprint PaliGemma-3B SigLIP-SO400M Open Image Page
RegionRAG (Li et al., 2025b) Preprint Qwen2.5-VL-3B ViT-BigG Open Image Element
IndustryRAG (Lim et al., 2025) EMNLP Industry Qwen2.5-VL-32B-Instruct ViT-BigG Open Image Page
COLMATE (Masry et al., 2025) EMNLP Industry PaliGemma-3B SigLIP-SO400M Open Image Page
LILaC (Yun et al., 2025) EMNLP $M$-Embed Open Image Element
HKRAG (Tong et al., 2025) Preprint Phi3V CLIP-ViT-L/14 Open Image Element
SLEUTH (Liu et al., 2025a) Preprint PaliGemma-3B SigLIP-SO400M Open Image Page
Snappy (Georgiou, 2025) Preprint PaliGemma-3B SigLIP-SO400M Open Image Element
@@ 3.1 Open and Closed Domain
Open-Domain Multimodal RAG. Open-domain multimodal RAG enhances an LLM’s knowledge in specialized domains by constructing retrieval databases from large collections of documents. Early approaches typically apply OCR to all documents to build text-based retrieval indices (Wang et al., 2022; Li et al., 2023; Chen et al., 2024c; Khattab and Zaharia, 2020), but this process is computationally expensive and inefficient. To improve scalability, recent methods such as DSE (Ma et al., 2024b) and ColPali (Faysse et al., 2024) leverage vision-language models (VLMs) to encode document pages directly into image embeddings, achieving significant efficiency gains. Despite these advances, most approaches still focus on reasoning within single documents and lack explicit mechanisms for integrating knowledge across sources. Addressing this limitation, M3DocRAG (Cho et al., 2024a) introduces approximate indexing to accelerate large-scale retrieval and establishes the benchmark M3DocVQA with over 3,000 documents, while VDocRAG (Tanaka et al., 2025) constructs the OpenDocVQA dataset and mitigates page-level information loss by compressing visual content into dense token representations aligned with text.
@@
Closed-Domain Multimodal RAG. Closed-domain multimodal RAG is designed for practical scenarios where MLLMs encounter difficulties with extremely long documents or videos. Current MLLMs remain constrained by limited context windows, and long-context processing often amplifies the risk of hallucination. To address this, closed-domain approaches retrieve only the most relevant segments (e.g., pages or frames) from a target document and provide them as input to the MLLM, thereby improving both efficiency and reliability. For single-document visual question answering (DocVQA), SV-RAG (Chen et al., 2024b) employs the MLLM itself as a multimodal retriever, with specialized adapters for page retrieval and evidence-based reasoning. FRAG (Huang et al., 2025a), by contrast, independently scores each frame or page, applies a Top-K selection to retain the most informative content, and then delegates answer generation to existing LMMs. CREAM (Zhang et al., 2024a) introduces a coarse-to-fine multimodal retrieval and attention-pooling integration framework, enabling effective cross-page reasoning and multi-page document comprehension for visual question answering. All approaches demonstrate that closed-domain multimodal RAG enables effective comprehension of long documents and videos without extending the model’s context length.
@@ 3.2 Retrieval Modality
Image-based Retrieval Modality. To handle visually rich documents, most existing methods represent each page as an image and encode it with VLMs, using the VLMs’ hidden states as page-level representations (see Figure 3 (a)). In parallel, the query is encoded, and page–query relevance is computed via similarity-based ranking (Ma et al., 2024b; Faysse et al., 2024; Yu et al., 2024; Chen et al., 2024b; Ma et al., 2025; Yu et al., 2025a). Building on image-based embeddings, MM-R5 (Xu et al., 2025a) introduces a reasoning-enhanced reranker that combines supervised fine-tuning and reinforcement learning to strengthen instruction following, elicit explicit reasoning chains, and leverage task-specific rewards for greater precision and interpretability. Complementing this direction, Light-ColPali (Ma et al., 2025) and HPC-ColPali (Bach, 2025) improve the efficiency of ColPali-style multi-vector retrieval by compressing patch-level representations, reducing memory and computation while largely preserving retrieval accuracy.
Image+Text based Retrieval Modality. Leveraging both image and text for retrieval mitigates the loss of fine-grained textual cues that arise when relying solely on page-level VLM encoders. The text channel is derived from OCR (Zhang et al., 2024a; Suri et al., 2025; Liu et al., 2025b; Wang et al., 2025b) or from summary annotations generated by large VLMs (Jain et al., 2025; Choi et al., 2025) (see Figure 2 (b)). VisDoMRAG (Suri et al., 2025) and HM-RAG (Liu et al., 2025b) adopt a dual-path pipeline: they retrieve and reason within each modality, then summarize and fuse the results into a single answer. By contrast, ViDoRAG (Wang et al., 2025b) and PREMIR (Choi et al., 2025) also retrieve per modality but merge candidates via a simple union before answer generation. Complementing these designs, SimpleDoc (Jain et al., 2025) uses a two-stage scheme for DocVQA: embedding-based candidate selection followed by re-ranking with VLM-generated page summaries, so that the summaries provide richer semantics for more precise evidence aggregation.
@@ 3.3 Retrieval Granularity
Recent multimodal RAG research demonstrates a clear evolution toward fine-grained, structure-aware evidence selection. VRAG-RL (Wang et al., 2025c) leverages reinforcement learning for region guidance, while MG-RAG (Xu et al., 2025b) and MMRAG-DocQA (Gong et al., 2025) enable multi-granularity retrieval via hierarchical indexing across pages and layouts. At the segment level, DocVQA-RAP (Yu et al., 2025a) ranks segments to suppress redundancy. Beyond segmentation, mKG-RAG (Yuan et al., 2025) aligns cross-modal entities via knowledge graphs, whereas PREMIR (Choi et al., 2025) matches queries against QA pairs for charts. Recent region-level methods like MARA (Wu et al., 2025a) and RegionRAG (Li et al., 2025b) introduce query-aligned representations and patch aggregation to reduce noise. Furthermore, HKRAG (Tong et al., 2025) captures fine-print knowledge via hybrid masking, and Snappy (Georgiou, 2025) achieves efficient localization by propagating patch-level similarity. Collectively, these approaches illustrate the shift toward increasingly fine-grained retrieval in document-heavy systems.
@@ 3.4 Hybrid Enhancements for Multimodal RAG
The main text focuses on integrating multimodal RAG with graph-based and agent-based methods. The Appendix G and  H extends this discussion to more advanced integrations, highlighting open challenges and future research directions.
@@
HM-RAG (Liu et al., 2025b) introduces a hierarchical multi-agent framework utilizing graph databases to capture structured relations, while mKG-RAG (Yuan et al., 2025) explicitly constructs multimodal knowledge graphs to align entities across vision and text. Building on such structured repositories, DB3Team-RAG (Xia et al., 2025) incorporates image-indexed graphs to handle complex ego-centric queries within domain-specific pipelines. Shifting focus to document topology, MoLoRAG (Wu et al., 2025b) leverages page graphs to encode logical connections for multi-page understanding. This structure-aware modeling is further refined by RECON (Wang and Chen, 2025), which builds a global graph linking intra-page visual relations with inter-page entity connections. Furthermore, LAD-RAG (Sourati et al., 2025) and LILaC (Yun et al., 2025) focus on layout-aware component graphs, employing dynamic traversal or late interaction to support multi-hop reasoning. Collectively, these methods highlight the pivotal role of graph structures as either external repositories or internal document representations in advancing reliable multimodal retrieval.
@@
Agent-based Multimodal RAG.
Agent-based multimodal RAG extends this paradigm by employing autonomous agents to orchestrate retrieval–generation interactions across modalities. These agents dynamically formulate queries, select retrieval strategies, and adaptively fuse information from text, images, tables, and other modalities (see Figure 5 (b)). Multi-agent collaboration further enables iterative reasoning, verification, and evidence refinement, improving the accuracy, reliability, and transparency of multimodal RAG systems.
ViDoRAG (Wang et al., 2025b) introduces an iterative agent workflow in which agents perform exploration, summarization, and reflection, improving multimodal retrieval and reasoning over visually rich documents. HM-RAG (Liu et al., 2025b) further extends this idea with a hierarchical multi-agent architecture, combining query decomposition, modality-specific parallel retrieval, and a decision agent that aggregates evidence through consistency voting and refinement. Adapting agentic RAG to the medical domain, Patho-AgenticRAG (Zhang et al., 2025c) enables task decomposition and multi-turn search to retrieve aligned text–image evidence from pathology textbooks while reducing diagnostic hallucinations. Along similar lines, HEAR (Chen et al., 2025a) and SLEUTH (Liu et al., 2025a) focus on improving long-document understanding by coupling VLM-based parsing with closed-loop or coarse-to-fine agent reasoning, allowing cross-modal inconsistencies to be corrected and salient evidence to be distilled into compact contexts. Overall, these approaches demonstrate how diverse agent designs enhance fine-grained retrieval and reasoning in multimodal RAG systems.